OpenStudy (keeli_grace):

Which is a solution of the equation? 2x + y = 5 A. (1, 2) B. 5 C. (2, –1) D. (3, –1) @Hero @ganeshie8 @phi @texaschic101

OpenStudy (anonymous):

D

OpenStudy (keeli_grace):

Thank you @Happyhappyhappy

OpenStudy (texaschic101):

do you know how she got the answer ?

OpenStudy (texaschic101):

she just subbed in her points into the equation to see if it satisfies it. (3,-1)...x = 3 and y = -1 2x + y = 5 2(3) + (-1) = 5 6 - 1 = 5 5 = 5 (correct) understand ? See how these points satisfy the equation ?

OpenStudy (keeli_grace):

Thank you @texaschic101 , I understand now. :)

OpenStudy (texaschic101):

good to hear :)

OpenStudy (anonymous):

A line and a point not on the line lie in exactly one plane.